OpenStudy (anonymous):

differentiate the following equation f(x)=((1/5)x)^5

OpenStudy (anonymous):

\[(\frac{ 1 }{ 5 }x)^5\]

sam (.sam.):

\[(\frac{1}{5}x)^5 = \frac{1}{5^5}x^5=\frac{1}{3125}x^5\] Using the power rule \[\frac{d}{dx}x^n=nx^{n-1}\] \[\frac{d}{dx}=\frac{5}{3125}x^{4}\]

sam (.sam.):

\[3125 \div 5 = 625\] \[\frac{d}{dx}=\frac{1}{625}x^{4}\]

OpenStudy (anonymous):

okay thanks :)